# Inheritune: Training Smaller Yet More Attentive Language Models

### v2

@@ 4.2 Ablations
5 Training without Data Repetition
(a) GPT2-Large
(b) GPT2-Medium
Figure 4: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using fewer layers. Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts until 80% of the training.
Are the gains we observe due to the Inheritune recipe merely a consequence of over-fitting due to data repetition?
To investigate this, we conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023a); Biderman et al. (2023). Moreover, we utilized a high-quality pre-training dataset, Fineweb_edu Penedo et al. (2024), which contains 100B tokens and has been deduplicated and filtered to ensure high data quality.
We train a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following Algorithm 1. Finally, we trained the 16-layer baseline variants from scratch for comparison. All these models are trained for 100K steps. The model configurations and training hyper-parameters can be found in supplementary Table 10.
As shown in Figure 4, our GPT-2 variants trained using $\mathsf{Inheritune}$ consistently perform on par with their full-size counterparts and outperform their same-size counterparts in terms of training loss. In LLM pre-training literature, where data is not repeated, training loss has been shown to be a reliable metric Touvron et al. (2023a, b). Furthermore, we performed zero-shot downstream evaluations using lm-evaluation-harness Gao et al. (2024) on a variety of tasks, including ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), Hellaswag Zellers et al. (2019), and PIQA Bisk et al. (2020). As shown in Table 4, on average, the models demonstrate superior performance than the baseline models trained from scratch.
@@ 5 Training without Data Repetition
Table 4: Models trained with Inheritune outperform both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
6 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied as a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation, i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggests that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to previous work, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023), which employs a stage-wise training strategy that uses weights from initial layers to initialize later layers, has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers-though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating a smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
7 Conclusion
In this study, we identified a structural flaw in the attention mechanism of deep decoder-style transformers, which are currently the de-facto architectural choice for LLMs. Specifically, we discovered that the attention matrices in many deeper layers tend to lose rank and converge into single-column matrices, resulting in what we term lazy layers. This phenomenon causes the model to lose attention and leads to structural inefficiencies in LLMs.
To avoid this issue, we propose a simple training recipe $\mathsf{Inheritune}$ where we inherit a few early blocks from a larger reference model and continue training and expanding the model with more blocks until it matches the performance of the reference model. We evaluated our method using 48-layer, 36-layer, and 24-layer GPT-2 xlarge, GPT-2 large, and GPT-2 medium reference models. Through this process, we successfully developed smaller models with 24, 18, and 16 layers without losing any performance compared to their larger counterparts while utilizing the OpenWebText-9B dataset. We observed thematically similar results with GPT-2 large and GPT-2 medium models trained on FineWeb_edu without data repetition. These findings offer a novel perspective on developing efficient small base language models from existing large pre-trained models, potentially democratizing LLM pre-training for a wider community.
Acknowledgments
We thank Nived Rajaraman, for his helpful feedback and discussion during this work.
Supplementary Materials
Contents
  • A: Supplementary Experiments
  • B: Additional Experiments on Low Data Regime
  • C: Implementation Details
Appendix A Supplementary Experiments
We provide additional training plots for our main results discussed in Section 4.1 as shown in Figure 5 and Figure 6. In Figure 5 (also refer Table 1 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 6 (also refer Table 2) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
Knowledge Distillation
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 2. Following up on our previous discussions in Figure 9 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers (refer Figure 9) compared to it’s larger counterpart trained form scratch. We performed rank analysis utilizing 24-layer GPT-2 medium and our 16-layer GPT-2 model trained using $\mathsf{Inheritune}$.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model trained from scratch, despite being smaller. Training GPT-2 xlarge, GPT-2 large and GPT-2 medium vanilla models from scratch and our variants with OpenWebText-9B for 100K steps.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 9: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers, indicating rank collapse. (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach in preserving expressive power.
@@ Appendix A Supplementary Experiments
Appendix B Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
@@
B.1 Main Results in Low Data Regime

### v3

@@ 4.2 Ablations
5 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied is a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). Recently He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023) employs a stage-wise training strategy that uses weights from initial layers to initialize later layers has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
6 Conclusion
In this paper, we identified a structural flaw in the attention mechanism of deep decoder-style LLMs, where many deeper layers tend to lose rank and converge into near rank-1 matrices. To address this, we propose $\mathsf{Inheritune}$, to train smaller models that inherits early blocks from a larger model and expands the architecture gradually, matching the performance of the reference model. We validated $\mathsf{Inheritune}$ on GPT-2 models of varying sizes, achieving efficient smaller models without performance loss on the OpenWebText and FineWeb_edu datasets.
Supplementary Materials
Contents
  • A: Extended Discussion on Attention Collapse
  • B: Baselines
  • C: Supplementary Experiments
  • D: Extended Discussion
  • E: Additional Experiments on Low Data Regime
  • F: Implementation Details
Appendix A Extended Discussion on Attention Collapse
A.1 Attention Mass Analysis
Dominant Single-Column Structure in Attention Matrices.
We further investigated the dominant structure of the rank-1 attention matrices and observed that, on an average, many of these matrices have their mass concentrated in a single column. This intrinsic structure can be viewed as a special case of rank-1 attention matrices. To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$.
Next we determine the minimal number of columns required to capture $\eta$ proportion of the total mass, formally computed as;
$m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq\eta\right\},$
Here $\eta\in(0,1)$ represents the cumulative column mass threshold. In this work, we set $\eta=0.90$. A lower value of $m^{*}$ implies a stronger concentration of the attention matrix mass within fewer columns, reinforcing the phenomenon attention collapse. This analysis provides additional quantitative evidence highlighting the reduced representational capability of attention matrices in deeper transformer layers, further supporting the identification of lazy layers.
In Figure 8, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models. For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer. We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$. We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion. As shown in Figure 9, we performed a mass analysis on contemporary billion-parameter LLaMA models (Geng & Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers. This provides concrete evidence that the phenomenon persists across a broad range of architectures at the billion-parameter scale.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 xLarge (1.5B) L=48, H=1600
Figure 8: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near single column matrices, which is a special case of near rank-1. We compute $\text{AvgMass}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set. Our mass analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models (L:layer, H:hidden size) reveals that attention matrices in many deeper layers collapse to single column matrices on an average.
(a) OpenLLaMA-3B
(b) OpenLLaMA-7B
(c) OpenLLaMA-13B
Figure 9: The overall mass of attention matrices in billion-scale LLMs, pre-trained on trillions of tokens, tends to concentrate in fewer columns. This phenomenon becomes increasingly pronounced as the model size grows. We computed attention matrices using 100 tokens from a random subset of RedPajama with 1B tokens. Next, we performed 100 runs and plotted the mean and standard deviation of the mass as a function of layers for our mass analysis, respectively. We followed the same procedure as discussed in Section 2. Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng & Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized. Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column. This observation aligns closely with our analysis in Figure 1.
A.2 Additional Analysis on Attention Rank Collapse.
To investigate the robustness of our analysis, we performed a rank analysis on GPT-2 models trained on OpenWebText (as shown in Figure 1) using a validation set from FineWeb-Edu—a new dataset on which the models were not trained (see Figure 10). It is clearly observed that attention consistently loses rank, often collapsing to single-column structures across all models, thereby reinforcing the robustness of our findings.
Next, we conducted a rank analysis on a contemporary LLaMA-3 model with 8B parameters. We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019). As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red. This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
A.3 Attention Pattern Visualization
To provide further evidence of lazy layers, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 12 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models computes causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over 3 runs.
In the 24-layer model trained from scratch (top row of Fig. 12), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes model more attentive and addresses attention degeneration, potentially leading to more efficient models in compact size (also refer Figure 17 and Figure 18). Additionally, in Figure 13, we present attention patterns from a 48-layer GPT-2 xLarge model, comparing attention maps from a potent (non-lazy) and a lazy layer.
(a) Rank analysis of GPT-2 Medium
(b) Rank analysis of GPT-2 Large
(c) Rank analysis of GPT-2 XLarge
(d) Matrix mass analysis of GPT-2 Medium
(e) Matrix mass analysis of GPT-2 Large
(f) Matrix mass analysis of GPT-2 XLarge
Figure 10: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations. All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the Fineweb_edu validation set, following the same procedure described in Figure 1. This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
Figure 11: Rank analysis of LLaMA-3 8B reveals that nearly half of the attention heads exhibit rank collapse. We analyze the LLaMA-3 8B model, which contains 32 heads per layer (32 $\times$ 32), using the rank metric defined in Section 2. The results are visualized as a heatmap of head index vs. layer index. Potent (non-collapsed) heads are shown in varying shades of green, where higher intensity indicates higher rank, while rank-collapsed heads (near rank-1) are highlighted in red. Approximately 50% of all attention heads exhibit rank collapse, indicating widespread degeneracy.
L4 H2
L7 H3
L20 H2
L22 H4
24 layer GPT2 medium model trained from scratch.
L4 H2
L7 H3
L11 H2
L15 H1
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
Figure 12: $\mathsf{Inheritune}$ preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected string, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s deeper layers.
L8 H16
L8 H17
L8 H18
L8 H19
L8 H20
A Potent (non-lazy) layer of a pre-trained GPT2 xLarge 48 layer model.
L30 H16
L30 H17
L30 H18
L30 H19
L30 H20
A lazy layer of a pre-trained GPT2 xLarge 48 layer model.
Figure 13: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 xLarge model with 48 layers. The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
Appendix B Baselines
We compare $\mathsf{Inheritune}$ against several baselines. While some baseline methods are illustrated using GPT-2 Large or medium (for the knowledge distillation baseline) as an example, the same methodology is consistently applied across all model variants.
Baselines trained from scratch (rand init.) :
Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various zeroshot model initialization and effcient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process for other models.
Stacking Gong et al. (2019); J. Reddi et al. (2023) is a model initialization and efficient (stage-wise) training recipe. We train a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17. Finally we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
Hybrid stacking: Hybrid stacking is stacking but utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights. We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to a 18-layer model by copying the weights to layers 0-17. We then trained this new 18-layer GPT-2 variant for 100K steps.
Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth. We copied the weights of the first half the attention heads (0-9) and MLPs of the GPT-2 large reference model into baseline GPT-2 variant with half the width but all layers.
Baselines trained with Knowledge Distillation
As a baseline, we first apply logit-based knowledge distillation Hinton et al. (2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly. For the second baseline, we use a DistillBERT-style approach Sanh et al. (2019), where the student model 0-11 layers are initialized with every alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher. Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model).
Appendix C Supplementary Experiments
We provide additional training data and plots for our main results discussed in Section 4.1 as shown in Table 3 and Figure 14. In Figure 5 (also refer Table 3 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 14 (also refer Table 1) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
Knowledge Distillation
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 12. Following up on our previous discussions in Figure 17 and Figure 18 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers compared to it’s larger counterpart trained form scratch. We performed rank analysis for Figure 17 utilizing vanilla 24-layer GPT-2 medium and our 16-layer GPT-2 variant trained using $\mathsf{Inheritune}$. Additionally, we performed rank analysis for Figure 18 with a vanilla 48-layer GPT2 xLarge and a 24-layer GPT2 xLarge variant trained using $\mathsf{Inheritune}$.
Models Layers Initialization Steps Pre-train Val loss ($\downarrow$)
GPT-2 Medium 24 rand init 100K 2.81
16 rand init 100K 2.86
12 Ours 100K 2.87
14 Ours 100K 2.84
Final Model $\longrightarrow$ 16 Ours 100K 2.81
GPT-2 Large 36 rand init 100K 2.85
18 rand init 100K 2.97
18 Ours 100K 2.80
GPT-2 xLarge 48 rand init 100K 2.65
24 rand init 100K 2.69
24 Ours 100K 2.64
Table 3: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts. All models are trained for 100K steps, and reported using validation loss on the OpenWebText validation set. Note: GPT-2 Large and xLarge use one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 17: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank accross all the 100 runs). (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
(a) Vanilla GPT2-xLarge 48 layers
(b) Our GPT2-xLarge 24 layer variant trained with $\mathsf{Inheritune}$
Figure 18: Rank collapse worsens for larger LLMs, $\mathsf{Inheritune}$ helps to mitigate rank collapse. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 48-layer GPT2 xLarge model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs). (b) Our 24-layer GPT2 xLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
Appendix D Extended Discussion
D.1 Discussion about Attention Sink
The term "attention sink" Xiao et al. (2024) refers to the phenomenon where the first token in a sequence receives disproportionately high attention scores compared to other tokens in the attention maps. While there is some connection with $\mathsf{Inheritune}$, as we have also observed that many attention matrices are not only rank-1 but also single-column (with all attention scores concentrated on the first token), this connection has not been explicitly established in Xiao et al. (2024) with respect to rank-1 behavior or poor training of later layers.
In contrast, as illustrated in Figure 1, we compute the maximum rank of all attention matrices within a layer. For instance, consider a layer where only 2 out of 5 attention heads exhibit attention sink behavior. This does not make the layer lazy, as attention is computed as a concatenation of activations across all heads. A lazy layer, however, has all 5 out of 5 attention heads fully degenerated, with their attention matrices being rank-1. We have provided evidence that such lazy layers are indicative of poorly trained layers.
Layers Initialization Avg max ranks Val Loss ($\downarrow$)
4 rand n/a 3.25
4 1-4 layers from vanilla GPT2 8.40 3.22
4 5-8 layers from vanilla GPT2 9.48 3.19
4 9-12 layers (lazy layers) from GPT2 1.22 3.23
Table 4: Impact of initialization strategies on GPT2-small variants. We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps. Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText. Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer. The training plots and rank analysis are provided in Figure 2.
@@ D.1 Discussion about Attention Sink
Table 5: Models trained with Inheritune outperforms both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
@@
Appendix E Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
@@
E.1 Main Results in Low Data Regime